# Soft-QMIX: Integrating Maximum Entropy For Monotonic Value Function Factorization

### v2

@@ 4.2 Soft Policy Iteration
$\pi_{jt}^{\text{new}}=\underset{\pi^{\prime}_{jt}\in\Pi}{\mathrm{arg\,min}}\ D_{KL}\left(\pi^{\prime}_{jt}(\cdot|s_{t})\middle\|\frac{\exp(Q_{tot}^{\pi_{\text{old}}}(s_{t},\cdot)/\alpha)}{Z^{\pi_{\text{old}}}(s_{t})}\right),$ (9)
[…]
In value-based methods, Equation 9 is intractable. Therefore, in the policy improvement step, we update the policy in this way:
@@
$f_{i}^{\text{new}}=\underset{f^{\prime}_{i}}{\mathrm{argmin}}\frac{1}{2}\left(\sum_{i}f^{\prime}_{i}(Q_{i}(o^{i}_{t},u^{i}_{t},s_{t}))-y_{t}\right)^{2},$ (10)
where $y_{t}$ can be any unbiased estimate of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u}_{t})\right]$. Here, we assume [^2] This does not necessarily hold in practical implementations, but the difference between these is a term in the loss function of our algorithm. For further discussion, please see the Appendix F. that $\mathcal{T}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u_{t}})$ is an unbiased estimate of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}\sum_{i}f_{i}(Q_{i}(s_{t},\mathbf{u_{t}}))\right]$. Under this assumption, we can replace $Q_{\theta}$ and $y_{t}$ in Theorem 4.2 with $\sum_{i}f_{i}(Q_{i})$ and $Q_{tot}$ respectively, and get that minimizing the L2 norm described in Equation 10 is equivalent to minimizing the KL divergence described in Equation 9 and minimizing the TD loss of value function at the same time. Therefore, in value-based methods, Equation 10 can be used in place of Equation 9.
@@
If $Q$-function is parameterized as $Q_{\theta}$, and $V_{\theta},\pi_{\theta}$ are defined as below:
$V_{\theta}(s)$ :-$\coloneq\alpha\log\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\exp\left(Q_{\theta}(s,\mathbf{u})/\alpha\right)\right],:$ (11)
$\pi_{jt_{\theta}}(\mathbf{u}|s)$ :-$\coloneq\exp\left(Q_{\theta}(s,\mathbf{u})-V_{\theta}(s)/\alpha\right).:$
and $y_{t}$ is an unbiased estimator of $\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\mathcal{T}^{\pi_{jt}}Q_{\theta}(s_{t},\mathbf{u_{t}})\right].$ Then, we have
$\nabla_{\theta}\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}$ $\left[\|Q_{\theta}(s,\mathbf{u})-y\|^{2}\right]$ (12)
$=\alpha\nabla_{\theta}D_{KL}\big{(}\pi_{jt}(\cdot|s)\|\frac{exp(y(\cdot,s)/\alpha)}{Z(s)}\big{)}+\nabla_{\theta}\mathbb{E}_{s,\mathbf{u}\sim\pi_{jt}}\left[\|V_{\theta}(s)-y\|^{2}\right]$
Proof. See Appendix E.
The subsequent theorem ensures that updating the policy using Equation 9 results in a monotonic increase of the Q-function under the IGI condition.
Lemma 4.3 (Joint Soft Policy Improvement).
Proof. See Appendix C.
Theorem 4.4 demonstrates that iteratively applying policy evaluation and policy improvement leads to a globally optimal policy under the IGI condition.
Theorem 4.4 (Joint Soft Policy Iteration).
By continually applying soft policy evaluation and soft policy improvement to any $\pi_{jt}\in\Pi$ converges to a policy $\pi^{*}$ such that $Q^{\pi^{*}}(s_{t},\mathbf{u}_{t})\geq Q^{\pi}(s_{t},\mathbf{u}_{t})$ for all $\pi_{jt}\in\Pi$ and $(s_{t},\mathbf{u}_{t})\in S\times\mathbf{U}$, assuming $|\mathbf{U}|<\infty$.
Proof. See Appendix C.
4.3 The Soft-QMIX Algorithm
Figure 2 illustrates the network architecture of the Soft-QMIX algorithm, which comprises four types of networks. The Mixer network is responsible for handling credit assignment and is parameterized along with its corresponding hypernetwork as $\text{Mixer}_{\theta}$. The $[g_{i}]_{i=1}^{n}$ and $[f_{i}]_{i=1}^{n}$ are $2n$ order-preserving transformations, with each $g_{i},f_{i}$ and their corresponding hypernetwork parameterized as $g_{i,\theta}$ and $f_{i,\theta}$. Agents 1 to $n$ share parameters, also referred to as $f_{i}^{\text{stage1}}$ in the previous section, and are parameterized as $f^{\text{stage1}}_{\theta}$. The updated rules for these parameters are described as follows.
Soft-QMIX aims to minimize two loss functions. $\text{Mixer}_{\theta}$, $g_{i,\theta}$ and $f^{\text{stage1}}_{\theta}$ are updated using the following loss function:
$L_{Q}(\theta)=\frac{1}{2}\left(Q_{tot_{\theta}}(s_{t},\mathbf{u}_{t})-\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}(s_{t},\mathbf{u}_{t})\right)^{2},$ (13)
where $Q_{tot_{\theta}}(\mathbf{u},s)=\text{Mixer}_{\theta}\big{(}g_{1,\theta}(f^{\text{stage1}}_{\theta}(o^{1},u^{1}),s),\ldots,$ $g_{n_{\theta}}(f^{\text{stage1}}_{\theta}(o^{n},u^{n}),s),s\big{)}$, and $\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}(s_{t},\mathbf{u}_{t})$ is defined in Equation 7. We utilize the technique of target Q networks to stabilize training, where the target Q network is parameterized as $\bar{\theta}$, with $\bar{\theta}$ being the exponentially weighted moving average of $\theta$. Subsequently, the parameter $\phi$ is updated using the following formula:
$L_{\pi}(\phi)=\frac{1}{2}\left(\sum_{i=1}^{n}f_{i,\phi}(Q_{i}(o^{i}_{t},u^{i}_{t}),s_{t})-Q_{tot}(s_{t},\mathbf{u}_{t})\right)^{2}.$ (14)
In this process, the function $f_{i,\phi}$ takes $Q_{i}(o^{i}_{t},u^{i}_{t})$ as input, but during backpropagation, gradients are truncated at $Q_{i}$. This truncation ensures that this loss function updates only $f_{i,\phi}$. After the update, the policy can be expressed in the following form:
$\pi_{i}(u^{i}|o^{i},s)=\text{softmax}(\frac{f_{i}\left(g_{i}\big{(}f^{stage1}(o^{i}),s\big{)},s\right)}{\alpha})$ (15)
It’s important to note that during execution, we only need to call $f^{stage1}$ and apply the argmax operation to it. In this paper, $\alpha$ is fixed during training. Finally, the algorithmic procedure of Soft-QMIX is shown in appendix A.
5 Experiments
In this section, we evaluate our algorithm using matrix games from prior works [27], demonstrating its ability to consistently reach the global optimum. Then, we compare Soft-QMIX with QMIX on MPE [18] and empirically demonstrate Soft-QMIX algorithm achieves state-of-the-art performance on the SMAC-v2 [3] benchmark. We compare our algorithm with four baseline algorithms. MAPPO [34] adapts the on-policy, single-agent PPO [26] algorithm for multi-agent use, utilizing a centralized value function. QMIX [23] is a value-based method employing a monotonic value decomposition to create multiple local deterministic policies from a shared value function. It uses an $\epsilon$-greedy strategy for exploration. FOP [36] implements a maximum entropy policy gradient approach with value factorization similar to QPLEX. HASAC [16] offers an alternative multi-agent soft actor-critic framework, facilitating sequential decision-making by agents in a centralized training with centralized execution context. We elaborate implementation details and hyperparameters in Appendix J.
5.1 Matrix Games
In this section, we demonstrate the effectiveness of our algorithm in the classic one-step matrix games. Table 1 presents a classic non-monotonic payoff matrix, often used in past papers [27, 31] to discuss the representational capability of critic networks. QMIX, due to its monotonicity constraint, often incurs moderate errors in fitting such payoff matrices. However, in practice, during the training process, we are primarily concerned with the value estimates of those joint actions that are frequently visited. For suboptimal joint actions with lower Q values, their errors do not significantly impact the estimated expected returns. As we use TD-$\lambda$ to estimate returns in our experiments, such errors can be further reduced. Conversely, we are more concerned with whether the joint actions with the highest probability of being selected have the globally maximal value function, as we aim for the highest returns when executing with a deterministic policy. For suboptimal joint actions, we also wish to visit them with a certain probability to ensure adequate exploration.
Table 1: Payoff matrix of the one-step matrix game.
@@ A B C
In the matrix game shown in Table 2, QMIX exhibits high estimation error due to its monotonicity constraints, and its deterministic policy lacks effective exploration, which leads it to learn only suboptimal joint actions with a reward of 00. FOP, with its network structure similar to QPLEX, can accurately estimate value functions, yet the corresponding policy learned tends to select suboptimal joint actions more frequently. Soft-QMIX has smaller estimation errors for optimal joint actions and larger errors for suboptimal ones. However, the joint action with the maximum Q value in Soft-QMIX matches the true optimal joint action, and the local policy has the highest probability of selecting the optimal joint action. Soft-QMIX possesses this capability because its decision-making is divided into two steps, and in the first step, it ensures the order of probabilities of each action being selected in the policy.
Table 2: (a)-(c) respectively demonstrate the results of Soft-QMIX, FOP, and QMIX in the one-step matrix game. The first row and column display the policies adopted by the row and column players after $10k$ steps of training, respectively. The 3$\times$3 matrix in the bottom right corner represents the $Q_{\text{tot}}$ ultimately learned by the algorithm. The results indicate that Soft-QMIX is the only method among all that consistently chooses the optimal action with the highest probability.
@@
(b) result of FOP
$\epsilon$ 1-$2\epsilon$ $\epsilon$
@@
(c) result of QMIX
5.2 Multi-Agent Particle Environment
In this section, we test our algorithm on three custumed Multi-Agent Particle Environment(MPE) [18] scenarios and show that Soft-QMIX do better exploration compare to QMIX and ablate through different choice of $\alpha$. Specifically, we use three different Spread map, where in each map, agents are tasked to navigate to all the landmarks without collide with each other. Harder map means longer initial distance between agents and landmarks. Details are available in Appendix I. Table 3 shows that Soft-QMIX significantly outperforms QMIX on the Hard and Medium map which require more exploration, but shows only a slight advantage on the Simple map. Additionally, we find that higher $\alpha$ values enhance exploration and performance. These findings support our methodological claim that integrating maximum entropy RL with MARL algorithms boosts exploration capabilities. For each scenario, we conduct four experiments across five seeds, reporting the average evaluation episode rewards after 10M environment steps and their standard deviation in our results table.
Table 3: Experiment result in three MPE scenarios, where Soft-QMIX has higher performance gain from QMIX on Hard and Medium map, which need more exploration. Higher $\alpha$ also lead to better exploration and higher performance gain.
eval episode QMIX Soft-QMIX(ours)
rewards $\alpha=0.1$ $\alpha=0.3$ $\alpha=1.0$
Spread(Hard) -171.7 $\pm$ 3.8 -173.7 $\pm$ 3.0 -170.8 $\pm$ 3.1 -168.7 $\pm$ 4.6 (+1.8%)
Spread(Medium) -138.4 $\pm$ 2.9 -139.0 $\pm$ 2.0 -137.4 $\pm$ 3.3 -134.9 $\pm$ 1.8 (+2.5%)
Spread(Simple) -111.9 $\pm$ 1.6 -116.2 $\pm$ 2.7 -114.4 $\pm$ 2.0 -111.4 $\pm$ 0.9 (+0.4%)
5.3 SMAC-v2
We evaluated Soft-QMIX on the SMACv2 [3] benchmark, including six scenarios: protoss_5_vs_5, terran_5_vs_5, zerg_5_vs_5, protoss_10_vs_10, terran_10_vs_10, and zerg_10_vs_10. The results, as shown in Figure 3, were obtained by training each scenario for $10$ million environmental steps, with each curve representing the average of 5 seeds, and HASAC with an average of 3 seeds. Soft-QMIX outperforms the baseline algorithms in all scenarios. MAPPO, being an on-policy algorithm, has lower sampling efficiency, resulting in lower win rates for models trained under 10M environmental steps. Training with MAPPO for 60M environmental steps can achieve win rates comparable to 10M QMIX. QMIX, lacking exploration, tends to converge to local optima, especially in the two Zerg scenarios. HASAC and FOP perform poorly in all the scenarios. Ultimately, Soft-QMIX improves win rates by $5\%$ to $24\%$ in all scenarios compared to baseline algorithms.
Figure 3: Experiment result in six SMAC-v2 scenarios, where Soft-QMIX achieved faster convergence speed and higher performance compared to baseline algorithms.

### v3

@@ 4.2 Training Scheme of OPT
$\underset{\pi^{\prime}_{jt}\in\Pi}{\mathrm{min}}\ D_{KL}\left(\pi^{\prime}_{jt}(\cdot|s_{t})\middle\|\frac{\exp(Q_{tot}^{\pi_{\text{old}}}(s_{t},\cdot)/\alpha)}{Z^{\pi_{\text{old}}}(s_{t})}\right),$ (3)
where $Z^{\pi_{\text{old}}}(s_{t})=\sum_{\mathbf{u}}\exp(Q_{tot}^{\pi_{\text{old}}}(s_{t},\mathbf{u})/\alpha)$ and $\alpha$ is the temperature in the softmax distribution.
To simplify and decentralize the policy improvement process, we propose replacing $Q_{tot}^{\pi_{\text{old}}}(s_{t},\cdot)$ with $\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t},\cdot),s_{t})$, which leads to:
@@
$\underset{\pi^{\prime}_{jt}\in\Pi}{\mathrm{min}}\ D_{KL}\left(\pi^{\prime}_{jt}(\cdot|s_{t})\middle\|\frac{\exp(\sum_{i}\frac{\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t},\cdot),s_{t})}{\alpha})}{Z^{\prime\pi_{\text{old}}}(s_{t})}\right),$ (4)
where $Z^{\prime\pi_{\text{old}}}(s_{t})$ is the normalization factor. The joint policy $\pi^{\prime}_{jt}$ is implemented as a set of independent local policies for decentralized execution, and we have the analytical solution for the local policy of agent $i$ as:
$\pi_{i}^{\text{new}}(a|o^{i}_{t})\propto\exp(\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t},a),s_{t})/\alpha).$ (5)
We can see that the policy is defined based on the local Q-value and utilizes the OPT to ensure that the relative order of policy logits aligns with the local Q-values.
There is a gap between $\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t},\cdot),s_{t})$ and $Q_{tot}^{\pi_{\text{old}}}(s_{t},\cdot)$. Also, we find that the policy improvement using Equation (4) no longer guarantees a monotonic increase in the global Q-value, unlike the approach based on Equation (3). Specifically, we have the following theorem:
@@
Denote the policy before the policy improvement as $\pi_{\text{old}}$ and the policy achieving the optimality of the improvement step as $\pi_{\text{new}}$. Updating the policy with Equation (3) ensures monotonic improvement in the global Q-value. That is, $Q^{\pi_{\text{old}}}_{tot}(s_{t},\mathbf{u_{t}})\leq Q_{tot}^{\pi_{\text{new}}}(s_{t},\mathbf{u_{t}}),\forall s_{t},\mathbf{u_{t}}.$
Algorithm 1 ME-IGM
1:Input: $\theta,\phi$
2:Initiate: ${\theta^{-}}\leftarrow\theta,D\leftarrow\emptyset$
3:for each iteration do
4:  for each environment step do
5:   get $\pi_{i}(u^{i}_{t}|o^{i}_{t},s_{t})$ via Eq. 10
6:   $u^{i}_{t}\sim\pi_{i}(u^{i}_{t}|o^{i}_{t},s_{t}),\forall i\in\{1,..,n\}$
7:   $s_{t+1}\sim\mathcal{P}(s_{t},\mathbf{u}_{t},\cdot)$
9:  end for
10:  for each gradient step do
11:   calculate $\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}$ via Eq. 12
12:   $\theta\leftarrow\theta-\lambda_{\theta}\nabla_{\theta}L(\theta)$
13:   $\phi\leftarrow\phi-\lambda_{\phi}\nabla_{\phi}L(\phi)$
14:   $\omega\leftarrow\omega-\lambda_{\omega}\nabla_{\omega}L(\omega)$
15:   ${\theta^{-}}\leftarrow\tau\theta+(1-\tau){\theta^{-}}$
16:  end for
17:end for
In contrast, updating the policy with Equation (4) only ensures that $Q^{\pi_{\text{old}}}_{tot}(s_{t},\mathbf{u_{t}})-\epsilon\leq Q_{tot}^{\pi_{\text{new}}}(s_{t},\mathbf{u_{t}}),\forall s_{t},\mathbf{u_{t}}$. The gap $\epsilon$ satisfies:
$\epsilon\leq\frac{\gamma}{1-\gamma}\mathbb{E}_{s_{t+1}\sim\mathcal{P}}\big{[}C\sqrt{2D_{KL}\big{(}\pi_{\text{old}}(\cdot|s_{t+1})|\pi_{\text{new}}(\cdot|s_{t+1})\big{)}}\big{]}$ (6)
where $\gamma$ is the discount factor, $\mathcal{P}(s_{t},\mathbf{u_{t}},\cdot)$ is the transition function and $C=\underset{\mathbf{u}_{t+1}}{\mathrm{max}}\ \Delta_{Q}(s_{t+1},\mathbf{u}_{t+1})\triangleq|Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}),s_{t+1})|$.
Proof is provided in the Appendix. This theorem implies that we must enforce $\epsilon\leq 0$ to guarantee monotonic policy improvement when using Equation (4) as the objective. This motivates us to minimize the factor $C$ and adopt the following objective to train the OPTs.
$\min_{{\text{OPT}_{1:n}}}\mathbb{E}_{s,\mathbf{u}\sim\rho_{\pi_{jt}^{new}}}[\Delta_{Q}(s,\mathbf{u})^{2}],$ (7)
$\pi_{jt}^{new}=(\pi_{1}^{new},\cdots,\pi_{n}^{new})$ and $\pi_{i}^{new}$ is defined in Eq. (5).
(a) Payoff Matrix
@@ A B C
(b) Result of ME-QMIX (Ours)
@@
(d) Result of QMIX
[…]
@@
[…]
5 The Overall Framework: ME-IGM
Figure 3 illustrates the framework of ME-IGM, which consists of three types of networks. Agents 1 to $n$ share a network $f_{\theta}$ for predicting local Q-values. The Mixer network, which aggregates local Q-values and manages the credit assignment among agents, is parameterized through its associated hyper-network as $\text{Mixer}_{\theta}$ [^4] Our algorithm is compatible with any Mixer network design that satisfies the IGM condition. For instance, during evaluation, we utilize the Mixer networks of QMIX and QPLEX, leading to two variants of our algorithm: ME-QMIX and ME-QPLEX.. $[\text{OPT}_{i,\phi}]_{i=1}^{n}$ are $n$ order-preserving transformations, and $\text{OPT}_{i,\phi}(f_{\theta}(o_{t}^{i}),s_{t})$ defines a local stochastic policy for agent $i$. The update rules for these networks are described as follows.
ME-IGM optimizes two loss functions to update $\theta$ and $\phi$ separately. $\text{Mixer}_{\theta}$ and $f_{\theta}$ are trained by minimizing $L(\theta)$:
$\mathbb{E}_{\tau_{t}}\left[\frac{1}{2}\left(Q_{tot_{\theta}}(s_{t},\mathbf{u}_{t})-\widehat{\mathcal{T}}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u}_{t})\right)^{2}\right]$ (8)
where $\tau_{t}=\{o_{t:t+1}^{1:n},u_{t:t+1}^{1:n},s_{t:t+1},r_{t}\}$, $Q_{tot_{\theta}}(\mathbf{u},s)=\text{Mixer}_{\theta}\big{(}f_{\theta}(o^{1},u^{1}),\ldots,f_{\theta}(o^{n},u^{n}),s\big{)}$, and the estimated Q-target $\widehat{\mathcal{T}}^{\pi_{jt}}Q_{tot}(s_{t},\mathbf{u}_{t})=r_{t}+\gamma Q_{tot,\theta^{-}}(s_{t+1},\mathbf{u}_{t+1})$. As a common practice, we utilize a target Q network $Q_{tot,\theta^{-}}$ to stabilize training, with $\theta^{-}$ being the exponentially weighted moving average of $\theta$. Notably, local actions $u_{t:t+1}^{1:n}$ are sampled from local stochastic policies $\pi_{i}$ (as defined in Equation (10)) during roll-out, rather than selecting the argmax of local Q-functions. This offers two advantages: (1) the softmax policy function enhances exploration, and (2) incorporating state $s$ in the sampling process allows for better utilization of global information during centralized training.
Further, to train $[\text{OPT}_{i,\phi}]_{i=1}^{n}$, we adopt a sample-based version of Equation (7), denoted as $L(\phi)$:
$\mathbb{E}_{\tau_{t}}\left[\left(\sum_{i=1}^{n}\text{OPT}_{i,\phi}(f_{\theta}(o^{i}_{t},u^{i}_{t}),s_{t})-Q_{tot,\theta}(s_{t},\mathbf{u}_{t})\right)^{2}\right]$ (9)
OPTs define the local policies for each agent as follows:
$\pi_{i}(u^{i}|o^{i},s)=\text{softmax}\left(\text{OPT}_{i,\phi}\left(f_{\theta}(o^{i}),s\right)/\alpha_{\omega}\right)$ (10)
$\alpha_{\omega}$ is the temperature and, similar to Haarnoja et al. (2018), can be updated by minimizing the following loss function:
$L(\omega)=\mathbb{E}_{\tau_{t}}\left[-\alpha_{\omega}\log\pi_{jt}(\mathbf{u}_{t}|s_{t})-\alpha_{\omega}\bar{\mathcal{H}}\right],$ (11)
where $\bar{\mathcal{H}}$ is a predefined target entropy. Note that the local policies (i.e., $\pi_{i}$) are only used to collect training roll-outs $\tau_{t}$. During execution, we would select the argmax of $f_{\theta}$, which does not require any centralized information (e.g., $s$), since OPTs preserve the order of the input vectors. The pseudo code of ME-IGM algorithm is presented in Algorithm 1.
6 Experiments
In this section, we first evaluate our method on a classic matrix game, demonstrating its capability to learn a globally optimal policy despite the non-monotonic nature of the overall payoff matrix. Next, we compare our algorithm against a set of baselines on the SMAC-v2 and Overcooked benchmark, highlighting its state-of-the-art performance in cooperative MARL. Moreover, we conduct ablation studies to evaluate the benefits of entropy-driven exploration, assess the necessity of incorporating the order-preserving transformation, and analyze the impact of key hyperparameters.
Figure 4: The mean return and standard deviation of QMIX, QPLEX, ME-QMIX , ME-QPLEX in Overcooked. It shows that ME-QMIX and ME-QPLEX achieve higher returns and exhibit faster convergence, compared to QMIX and QPLEX which do not adopt maximum entropy.
MARL ME-IGM (Ours) IL+MARL
Task Scenario MAPPO IPPO QMIX QPLEX HASAC FOP ME-QMIX ME-QPLEX IMAX-PPO InQ
Protoss 5_vs_5 58.0 54.6 70.2 53.3 22.9 37.8 75.9$\pm$3.0 69.8$\pm$1.8 68.1 78.7
10_vs_10 58.3 58.0 69.0 53.7 12.1 8.6 78.5$\pm$2.7 78.7$\pm$3.4 59.6 79.8
10_vs_11 18.2 20.3 42.5 22.8 5.7 0.4 50.4$\pm$2.9 44.9$\pm$3.2 21.3 48.7
20_vs_20 38.1 44.5 69.7 27.2 - 2.0 73.9$\pm$3.0 38.3$\pm$4.0 76.3 80.6
20_vs_23 5.1 4.1 16.5 4.8 - 0.3 23.4$\pm$1.0 18.7$\pm$1.3 11.8 24.2
Terran 5_vs_5 52.0 56.2 58.4 70.0 37.9 57.1 70.2$\pm$2.6 70.0$\pm$4.8 53.3 69.9
10_vs_10 58.1 57.3 65.8 66.1 20.4 15.8 72.6$\pm$1.8 69.4$\pm$0.2 58.4 72.2
10_vs_11 28.6 31.0 39.4 41.4 9.1 7.5 49.6$\pm$1.6 51.0$\pm$0.4 28.4 53.9
20_vs_20 52.8 49.6 57.6 23.9 - 0.2 59.3$\pm$1.3 45.1$\pm$7.2 35.9 65.4
20_vs_23 11.2 10.0 10.0 7.0 - 0.0 18.7$\pm$0.7 17.4$\pm$2.9 4.7 17.7
Zerg 5_vs_5 41.0 37.2 37.2 47.8 29.1 35.2 50.8$\pm$4.6 52$\pm$3.9 48.6 55.0
10_vs_10 39.1 49.4 40.8 41.6 17.9 8.0 49.1$\pm$0.6 42.9$\pm$1.8 50.6 57.6
10_vs_11 31.2 26.0 28.0 31.1 14.0 1.5 32.6$\pm$1.7 45.3$\pm$0.6 34.8 41.5
20_vs_20 31.9 31.2 30.4 15.8 - 0.2 34.3$\pm$5.2 42.1$\pm$5.1 26.7 43.3
20_vs_23 15.8 8.3 10.1 6.7 - 0.2 14.8$\pm$6.4 17.5$\pm$3.2 8.2 21.3
Table 2: We record the average win rates of different algorithms on SMAC-v2. The results demonstrate that our method, ME-IGM, outperforms all previous baseline algorithms and achieves performance comparable to IL+MARL methods, despite not having access to expert opponent demonstrations or opponent modeling.
6.1 Evaluation on Matrix Games
We first demonstrate the effectiveness of our algorithm in a classic one-step matrix game, shown as Table 1. It is a non-monotonic payoff matrix commonly used in previous studies (Son et al. 2019; Wang et al. 2020). We evaluate ME-QMIX, a variant of ME-IGM built upon QMIX, by comparing its performance against two state-of-the-art MARL methods: QMIX (Rashid et al. 2020), a value decomposition MARL method, and FOP (Zhang et al. 2021), a Maximum Entropy MARL method. We are concerned with whether the joint action with the highest selection probability corresponds to the globally maximal Q value, ensuring optimal returns when executing a deterministic policy.
As shown in Table 1, QMIX suffers from high estimation error in payoff values of this non-monotonic game, due to its monotonicity constraints (as introduced in Section 2.3). Its deterministic policy converges to a suboptimal joint action (B, B) with a reward of 0. FOP, which employs a network architecture similar to QPLEX, is capable of accurately estimating value functions (i.e., the payoff). However, its policy tends to select suboptimal joint actions more frequently due to the misalignment between local policies and the maximum global Q-values, as discussed in Section 3. ME-QMIX exhibits smaller estimation errors for the payoff of optimal joint actions (which is our primary concern) while having larger errors for suboptimal ones. Crucially, the joint action with the maximum Q-value in ME-QMIX aligns with the true optimal joint action, and the local policy assigns the highest probability to selecting this optimal joint action, avoiding the misalignment issue through the use of OPTs.
6.2 Evaluation on Overcooked
To demonstrate the advantage of employing the principle of maximum entropy, we compare ME-IGM with two leading value decomposition MARL methods (which do not involve maximum entropy): QMIX and QPLEX, on the Overcooked benchmark (Carroll et al. 2019) – a widely used coordination-focused multi-agent environment. In particular, we evaluate them in two representative layouts: cramped room and asymmetric advantages. These scenarios pose distinct challenges, including close collaboration under spatial constraints and asymmetric advantages in task division, and require coordinated exploration. For each scenario, we conduct repeated experiments with different random seeds to get the training curves of 10 million environment steps. As shown in Figure 4, ME-QMIX and ME-QPLEX exhibit significantly faster convergence and achieve higher performance levels in the early stages of training compared to their counterparts, QMIX and QPLEX.
6.3 Evaluation on SMAC-v2
We further conduct a comprehensive evaluation of ME-IGM on the SMACv2 (Ellis et al. 2022) benchmark across 15 scenarios, comparing its performance against 8 baselines. The results, as shown in Table 2, are obtained by training each algorithm on each scenario for 101010 millions of environment steps. Each value in the table represents the average performance over three runs with different random seeds. ME-IGM outperforms the baseline algorithms across all scenarios by a significant margin. MAPPO and IPPO, as on-policy algorithms, have lower sampling efficiency, resulting in lower win rates within  [^10] M environment steps. QMIX and QPLEX, lacking exploration, tend to converge to local optimum. HASAC (Liu et al. 2023) and FOP perform poorly across all scenarios. Surprisingly, ME-IGM achieves performance on par with (or even surpassing) IL+MARL methods, despite not relying on expert opponent demonstrations or opponent modeling. The results of MAPPO, IPPO, QMIX, QPLEX, IMAX-PPO, InQ are taken from (Bui, Mai, and Nguyen 2024). A detailed comparison with actor-critic baselines is presented in the Appendix.